OpenStudy (anonymous):

What is the limit as x -> 0 of the function, f(x)= 1-cosx / sinx

OpenStudy (anonymous):

Do you know L'Hopital's rule? If not, there's a way to do it without it.

OpenStudy (anonymous):

\[\lim_{x\to0}\frac{1-\cos x}{\sin x}=\frac{0}{0}\] Apply L'Hopital's rule: \[\lim_{x\to0}\frac{1-\cos x}{\sin x}=\lim_{x\to0}\frac{\sin x}{\cos x}\] If you don't know L'Hopital's rule, you can use the fact that \(\sin x\approx x\) for \(x\) near 0, so you have \[\lim_{x\to0}\frac{1-\cos x}{\sin x}=\lim_{x\to0}\frac{1-\cos x}{x}\] which is a limit you may have already been exposed to.